OpenStudy (anonymous):

Hey, can anyone show me how to solve this step by step i'm really confused solve this polynomial equation below x^3-7x^2-x+7=0

OpenStudy (jiteshmeghwal9):

x^2(x-7)-1(x-7)=0 (x^2-1)(x-7)=0 x=1 & x=7

OpenStudy (anonymous):

oh so all i had to do was factor it?